OpenStudy (anonymous):

Let A =[ up 1,3,2; middle: 2,0, -1; bottom 1,1,0] then det A =2 and A^-1 ?

OpenStudy (anonymous):

my cofactor is \[\left[\begin{matrix}1 & 2&-3 \\ -1 & -2&2\\2&2&-6\end{matrix}\right]\]

OpenStudy (anonymous):

I mean transpose already, just plug 1/2 in front to get A^-

OpenStudy (e.mccormick):

Wait. Cofacor or transposed matrix of minors there?

OpenStudy (anonymous):

yap

OpenStudy (e.mccormick):

\[M_{3,2}=\left| \begin{matrix}1&2\\2&-1\end{matrix} \right|\implies M_{3,2}=-5\]Does that correct your issues?

OpenStudy (anonymous):

yes, thanks . that's it. I know my mistake

OpenStudy (e.mccormick):

Kk. Have fun!